# Multi-View Graph Learning with Graph-Tuple

### v1

@@ 1 Introduction
Graph neural networks (GNNs) have demonstrated remarkable success in learning from structured data [1], achieving state-of-the-art results across diverse fields such as social network analysis, recommendation systems, drug discovery, and materials science [2, 3, 4, 5, 6]. The power of GNNs stems from their ability to learn rich representations of nodes and entire graphs by iteratively passing and aggregating messages over a relational structure [5, 7]. This core mechanism endows them with a strong relational inductive bias [8]: the inherent assumption that an object’s properties are determined by its connections and local context. This bias is precisely why GNNs are so effective for tasks on graph-structured data [2].
@@
Computationally, GNNs typically scale with the number of graph edges, making them efficient for sparse graphs. However, this efficiency degrades as the graphs become denser. This poses a particular challenge on large, dense graphs such as fully-connected distance graphs derived from point clouds. To make GNN training efficient on dense graphs, a common way is to sparsify it by applying similarity thresholding or distance pruning [9]. Yet, this often reduces information and results in graph representations based on a single fixed scale of interaction. For example, a high threshold applied to a molecule’s Coulomb matrix retains only strong chemical bonds at the expense of losing important weaker connections. Conversely, a low threshold preserves these weaker connections, but also introduces significant noise. Alternatively, invariant feature models [10] reduce the computational cost to linear in the number of points, by exploiting the low-rank structure of the point cloud and allowing exact reconstruction of the full adjacency matrix from a small submatrix and anchor points.
@@
A complementary line of works avoids graph sparsification and low-rank assumptions, by decomposing a single dense graph into multiple sparser graphs and then learning them in parallel, such as multi-view methods  [11, 12, 13] and heterogeneous GNNs [14, 15]. These approaches preserve diverse interaction ranges while being computationally tractable, yet are typically designed for heterogeneous graphs with multiple node types or edges types, not directly applicable on homogeneous graphs with continuous edge features.
@@
To tackle these challenges, we propose a multi-view graph representation that captures both fine-grained and contextual interactions. Instead of a single graph, we construct a graph tuple over the same nodes by explicitly partitioning edges according to interaction strength (e.g., distance or Coulomb energy): a strong-connection graph retaining the strongest local interactions and a complementary weak-connection graph providing broader context. Inspired by the theoretical insights of the GtNN framework [16], we then explicitly integrate multiple message-passing operations in a single layer: intra-scale operations (within each graph view) and, crucially, inter-scale operations that model the distinct operator orderings (across different graph views). This yields an interpretable and physically grounded mechanism that links local topology to global effects. We prove that under mild assumptions, this heterogeneous message-passing architecture is more expressive than single-graph models and guaranteed to achieve a lower or equal oracle prediction risk.
@@
We instantiate our framework on two scientific domains. For molecular property prediction on the QM7b dataset, we develop GINE-Gt, a specialized architecture that uses the powerful Graph Isomorphism Network with Edge Features(GINE) [17] as the backbone. Second, for cosmological parameter inference from point cloud data, we develop EGNN-Gt based on Equivariant Graph Neural Network (EGNN) [18], a powerful GNN architecture that guarantees equivariance to rotations, translations, and reflections.
@@ 2 Related Work
Heterogeneous graph learning. Early heterogeneous GNNs such as R-GCN [14] and HAN [15] demonstrated the benefit of relation-specific message-passing design, but they assume pre-defined, discrete node and edge types (e.g., knowledge graphs or bibliographic networks). To go beyond pre-defined relations, Graph Transformer Networks (GTN) [11] proposed to softly select relation-specific adjacency matrices and then generate new graphs by their matrix products. We generalize this heterogeneous graph learning paradigm to homogeneous graphs, by inducing different relations via partitioning continuous edge features (e.g., physical distances or chemical interactions).
Multi-view graph learning. A recent line of work constructs multiple relational views from a single graph, motivated from self-supervising learning (e.g., contrastive multi-view learning [12]) or community detection (e.g., variational edge partition model [19]). In contrast to these prior works, we are motivated to construct multiple views based on a physical measure of interaction strength from scientific applications (e.g., the Coulomb matrix in the molecular domain, and the Euclidean distance matrix in the cosmological applications).
Multi-scale GNNs. Our approach is architecturally most related to multi-scale GNNs that learn a hierarchical representation from a graph, such as FraGAT [20] designed for molecular property prediction,and MultiScale MeshGraphNets [21] for physics simulation. But these methods typically build multiple graphs on different node sets (e.g., atoms vs. fragments and fine vs. coarse mesh), which requires non-trivial cross-level alignment and transfer operators. In contrast, our framework defines multiple graphs over the same node set by partitioning a continuous interaction strength, enabling simple and interpretable within- and between-graph message passing.
@@ 3.1 Graph Isomorphism Network with Edge Features (GINE)
GINE [17] extends Graph Isomorphism Network (GIN) by incorporating edge features into its message passing procedure:
@@ 3.2 Equivariant Graph Neural Networks (EGNN)
Equivariant Graph Neural Networks (EGNNs) [18] incorporate geometric information by endowing each node with Euclidean coordinates $x_{i}\in\mathbb{R}^{d}$. They jointly update node features and coordinates in a way that is equivariant to node permutations and Euclidean isometries, i.e., translations, rotations (and reflection). An EGNN convolution layer (EGCL) is defined as follows:
@@ 4 Method
Our work introduces a multi-view graph-tuple framework for learning from complex relational systems where interactions occur at different scales. The core principle is to model these interactions and learn how information flows both within and between these scales. This is achieved by decomposing a graph $\mathcal{G}$ into a graph tuple, $(\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{k})$, that provides distinct yet complementary views of the interaction space, representing fine-grained local structures and broader contextual relationships. This graph-tuple framework can be instantiated using different graph neural network backbones. We present two such instantiations: an edge-aware model for attributed graphs, GINE-Gt based on [17], and an equivariant extension for geometric data, EGNN-Gt [18].
@@ 4.2 Heterogeneous Message-Passing Architecture
Figure 1 provides an overview of our framework. We note that if the original edge set $\mathcal{E}$ contains typed edges, it can be naturally partitioned into homogeneous edge sets, one for each type, as done in R-GCN [14]. Our framework extends R-GCN from heterogeneous graphs to homogeneous graphs, replacing discrete edge types with partitions induced by continuous edge features. The architectures proposed below also extend the Graph Tuple Neural Network framework from [16].
@@ 4.2.1 GINE-Gt
For general graphs with node features and edge attributes, we implement our framework using the GINE layer [17]. The intra-scale message-passing for $i=1,\ldots,k$ are computed as
@@ 4.2.2 EGNN-Gt
For geometric data where node features represent point positions in $\mathbb{R}^{d}$, we provide an $E(d)$-equivariant implementation of our framework using the EGCL layer [18]. The overall feature update follows Eq. 7, while the position feature update is analogously defined as:
@@ 6 Molecular Property Prediction
We consider the QM7b benchmark [23, 24], which contains 7 2117\,211 molecules with $14$ regression targets. Each molecule is encoded by a $n\times n$ Coulomb matrix $X$ whose entries depend only on nuclear charges $Z_{i}\!\in\!\mathbb{R}$ and 3D coordinates $R_{i}\!\in\!\mathbb{R}^{3}$:
@@
Then we build the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $\mathcal{V}=\{1,\dots,n\}$ and $\mathcal{E}=\{(i,j)\mid i\neq j\}$, i.e. all atom pairs are connected while self–loops $(i,i)$ are removed.
@@ 6.2 Results and Analysis
Table 1 presents the performance comparison of our GINE-Gt model against multiple baselines. These include a series of single-graph GINE-$c$ models as well as several non-graph-based methods: Kernel Ridge Regression (KRR), Deep Tensor Neural Network (DTNN) [25], and the state-of-the-art invariant feature model, DS-CI+ [10]. Results for KRR and DTNN are taken from prior work [22].
@@ 7 Cosmological Parameter Prediction
We evaluate our EGNN-Gt model on the CosmoBench benchmark [9], specifically using the CAMELS (TNG) and the CAMELS-SAM cosmological point cloud datasets. Each sample in these datasets is a cosmological simulation cloud where nodes represent dark matter halos or galaxies. The model input is the matrix of 3D halo (galaxy) positions representing their present-day configuration, $X\in\mathbb{R}^{N\times 3}$, where $N$ is the number of halos in the point cloud.
@@ 7.1 Graph Construction
Insights from prior work [9] and our own preliminary experiments suggest that informative halo interactions typically occur within a radius range approximately from 0 to 10 Mpc/h. To establish strong baselines, we evaluated two types of single-graph models. First, we form a series of strong-connection baseline, EGNN-$c_{1}$ by connecting halos within a cutoff radius $c_{1}$. Second, we create dense-graph baselines, EGNN-$c_{2}$ using a larger radius $c_{2}>c_{1}$. For our EGNN-Gt model, we implement the multi-view approach by partitioning the interaction space with two radii: a strong-connection radius $c_{1}$ and a weak-connection radius $c_{2}$. To explore the benefits of multi-view processing while maintaining a simple and interpretable relationship between the two scales, we enforce a fixed ratio by setting $c_{2}=2c_{1}$ in all our experiments. This allows the model to simultaneously capture immediate local neighborhoods ($\mathcal{G}_{1}$) and broader, second-order contextual regions ($\mathcal{G}_{2}$). The specific ranges for the systematic search over $c_{1}$ for each task, along with full implementation details, are provided in Appendix B.2.
@@ B.1 Molecular Property Prediction
Since the QM7b dataset is feature-scarce, we first construct node and edge features from the molecule’s Coulomb matrix, $X$. Following prior work [10], we derive initial features directly from the Coulomb matrix entries. Specifically, we apply a "binary expansion" technique to expand the scalar diagonal entries ($X_{ii}$) and off-diagonal entries ($X_{ij}$) into 100-dimensional vectors, which serve as the initial node and edge features, respectively. These raw scalar values are then projected into a 100-dimensional hidden space by learnable encoders.
@@
For training, all models use a batch size of 128. We train the models by minimizing the L1 Loss (Mean Absolute Error) using the Adam optimizer [26] with an initial learning rate of $5\times 10^{-3}$ and weight decay of $10^{-5}$. A cosine–plateau scheduler reduces the learning rate by a factor of 0.8 after five epochs without validation improvement (minimum $10^{-5}$). Early stopping is triggered after 20 idle epochs or when the run reaches a maximum of 1000 epochs.

### v4

@@ 1 Introduction
Graph neural networks (GNNs) have demonstrated remarkable success in learning from structured data [bronstein2017geometric], achieving state-of-the-art results across diverse fields such as social network analysis, recommendation systems, drug discovery, and materials science [kipf2016semi, hamilton2017inductive, he2020lightgcn, gilmer2017neural, xie2018crystal]. The power of GNNs stems from their ability to learn rich representations of nodes and entire graphs by iteratively passing and aggregating messages over a relational structure [gilmer2017neural, xu2018powerful]. This core mechanism endows them with a strong relational inductive bias [battaglia2018relational]: the inherent assumption that an object’s properties are determined by its connections and local context. This bias is precisely why GNNs are so effective for tasks on graph-structured data [kipf2016semi].
@@
Computationally, GNNs typically scale with the number of graph edges, making them efficient for sparse graphs. However, this efficiency degrades as the graphs become denser. This poses a particular challenge on large, dense graphs such as fully-connected distance graphs derived from point clouds. To make GNN training efficient on dense graphs, a common way is to sparsify it by applying similarity thresholding or distance pruning [huang2025cosmobench]. Yet, this often reduces information and results in graph representations based on a single fixed scale of interaction. For example, a high threshold applied to a molecule’s Coulomb matrix retains only strong chemical bonds at the expense of losing important weaker connections. Conversely, a low threshold preserves these weaker connections, but also introduces significant noise. Alternatively, invariant feature models [blum2024learning] reduce the computational cost to linear in the number of points, by exploiting the low-rank structure of the point cloud and allowing exact reconstruction of the full adjacency matrix from a small submatrix and anchor points.
@@
A complementary line of works avoids graph sparsification and low-rank assumptions, by decomposing a single dense graph into multiple sparser graphs and then learning them in parallel, such as multi-view methods  [yun2019graph, hassani2020contrastive, park2020unsupervised] and heterogeneous GNNs [schlichtkrull2018modeling, wang2019heterogeneous]. These approaches preserve diverse interaction ranges while being computationally tractable, yet are typically designed for heterogeneous graphs with multiple node types or edges types, not directly applicable on homogeneous graphs with continuous edge features.
@@
To tackle these challenges, we propose a multi-view graph representation that captures both fine-grained and contextual interactions. Instead of a single graph, we construct a graph tuple over the same nodes by explicitly partitioning edges according to interaction strength (e.g., distance or Coulomb energy): a strong-connection graph retaining the strongest local interactions and a complementary weak-connection graph providing broader context. Inspired by the theoretical insights of the GtNN framework [velasco2024graph], we then explicitly integrate multiple message-passing operations in a single layer: intra-scale operations (within each graph view) and, crucially, inter-scale operations that model the distinct operator orderings (across different graph views). This yields an interpretable and physically grounded mechanism that links local topology to global effects. We prove that under mild assumptions, this heterogeneous message-passing architecture is more expressive than single-graph models and guaranteed to achieve a lower or equal oracle prediction risk.
@@
We instantiate our framework on two scientific domains. For molecular property prediction on the QM7b dataset, we develop GINE-Gt, a specialized architecture that uses the powerful Graph Isomorphism Network with Edge Features(GINE) [hu2019strategies] as the backbone. Second, for cosmological parameter inference from point cloud data, we develop EGNN-Gt based on Equivariant Graph Neural Network (EGNN) [satorras2021n], a powerful GNN architecture that guarantees equivariance to rotations, translations, and reflections.
@@ 2 Related Work
Heterogeneous graph learning. Early heterogeneous GNNs such as R-GCN [schlichtkrull2018modeling] and HAN [wang2019heterogeneous] demonstrated the benefit of relation-specific message-passing design, but they assume pre-defined, discrete node and edge types (e.g., knowledge graphs or bibliographic networks). To go beyond pre-defined relations, Graph Transformer Networks (GTN) [yun2019graph] proposed to softly select relation-specific adjacency matrices and then generate new graphs by their matrix products. We generalize this heterogeneous graph learning paradigm to homogeneous graphs, by inducing different relations via partitioning continuous edge features (e.g., physical distances or chemical interactions).
Multi-view graph learning. A recent line of work constructs multiple relational views from a single graph, motivated from self-supervising learning (e.g., contrastive multi-view learning [hassani2020contrastive]) or community detection (e.g., variational edge partition model [he2022variational]). In contrast to these prior works, we are motivated to construct multiple views based on a physical measure of interaction strength from scientific applications (e.g., the Coulomb matrix in the molecular domain, and the Euclidean distance matrix in the cosmological applications).
Multi-scale GNNs. Our approach is architecturally most related to multi-scale GNNs that learn a hierarchical representation from a graph, such as FraGAT [zhang2021fragat] designed for molecular property prediction,and MultiScale MeshGraphNets [fortunato2022multiscale] for physics simulation. But these methods typically build multiple graphs on different node sets (e.g., atoms vs. fragments and fine vs. coarse mesh), which requires non-trivial cross-level alignment and transfer operators. In contrast, our framework defines multiple graphs over the same node set by partitioning a continuous interaction strength, enabling simple and interpretable within- and between-graph message passing.
@@ 3.1 Graph Isomorphism Network with Edge Features (GINE)
GINE [hu2019strategies] extends Graph Isomorphism Network (GIN) by incorporating edge features into its message passing procedure:
@@ 3.2 Equivariant Graph Neural Networks (EGNN)
Equivariant Graph Neural Networks (EGNNs) [satorras2021n] incorporate geometric information by endowing each node with Euclidean coordinates $x_{i}\in\mathbb{R}^{d}$. They jointly update node features and coordinates in a way that is equivariant to node permutations and Euclidean isometries, i.e., translations, rotations (and reflection). An EGNN convolution layer (EGCL) is defined as follows:
@@ 4 Method
Our work introduces a multi-view graph-tuple framework for learning from complex relational systems where interactions occur at different scales. The core principle is to model these interactions and learn how information flows both within and between these scales. This is achieved by decomposing a graph $\mathcal{G}$ into a graph tuple, $(\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{k})$, that provides distinct yet complementary views of the interaction space, representing fine-grained local structures and broader contextual relationships. This graph-tuple framework can be instantiated using different graph neural network backbones. We present two such instantiations: an edge-aware model for attributed graphs, GINE-Gt based on [hu2019strategies], and an equivariant extension for geometric data, EGNN-Gt [satorras2021n].
@@ 4.2 Heterogeneous Message-Passing Architecture
Figure 1 provides an overview of our framework. We note that if the original edge set $\mathcal{E}$ contains typed edges, it can be naturally partitioned into homogeneous edge sets, one for each type, as done in R-GCN [schlichtkrull2018modeling]. Our framework extends R-GCN from heterogeneous graphs to homogeneous graphs, replacing discrete edge types with partitions induced by continuous edge features. The architectures proposed below also extend the Graph Tuple Neural Network framework from [velasco2024graph].
@@ 4.2.1 GINE-Gt
For general graphs with node features and edge attributes, we implement our framework using the GINE layer [hu2019strategies]. The intra-scale message-passing for $i=1,\ldots,k$ are computed as
@@ 4.2.2 EGNN-Gt
For geometric data where node features represent point positions in $\mathbb{R}^{d}$, we provide an $E(d)$-equivariant implementation of our framework using the EGCL layer [satorras2021n]. The overall feature update follows Eq. 7, while the position feature update is analogously defined as:
@@ 6 Molecular Property Prediction
We consider the QM7b benchmark [blum2009970, montavon2013machine], which contains 7 2117\,211 molecules with $14$ regression targets. Each molecule is encoded by a $n\times n$ Coulomb matrix $X$ whose entries depend only on nuclear charges $Z_{i}\!\in\!\mathbb{R}$ and 3D coordinates $R_{i}\!\in\!\mathbb{R}^{3}$:
@@
Then we build the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $\mathcal{V}=\{1,\dots,n\}$ and $\mathcal{E}=\{(i,j)\mid i\neq j\}$, i.e. all atom pairs are connected while self–loops $(i,i)$ are removed. Notably, since its off-diagonal entries are computed from inter-atomic distances, the Coulomb matrix is invariant to rotations and translations (i.e., E(3)-invariant) by construction.
@@ 6.2 Results and Analysis
Table 1 presents the performance comparison of our GINE-Gt model against multiple baselines. These include a series of single-graph GINE-$c$ models as well as several non-graph-based methods: Kernel Ridge Regression (KRR), Deep Tensor Neural Network (DTNN) [schutt2017quantum], and the state-of-the-art invariant feature model, DS-CI+ [blum2024learning]. Results for KRR and DTNN are taken from prior work [wu2018moleculenet].
@@ 7 Cosmological Parameter Prediction
We evaluate our EGNN-Gt model on the CosmoBench benchmark [huang2025cosmobench], specifically using the CAMELS (TNG) and the CAMELS-SAM cosmological point cloud datasets. Each sample in these datasets is a cosmological simulation cloud where nodes represent dark matter halos or galaxies. The model input is the matrix of 3D halo (galaxy) positions representing their present-day configuration, $X\in\mathbb{R}^{N\times 3}$, where $N$ is the number of halos in the point cloud.
@@ 7.1 Graph Construction
Insights from prior work [huang2025cosmobench] and our own preliminary experiments suggest that informative halo interactions typically occur within a radius range approximately from 0 to 10 Mpc/h. To establish strong baselines, we evaluated two types of single-graph models. First, we form a series of strong-connection baseline, EGNN-$c_{1}$ by connecting halos within a cutoff radius $c_{1}$. Second, we create dense-graph baselines, EGNN-$c_{2}$ using a larger radius $c_{2}>c_{1}$. For our EGNN-Gt model, we implement the multi-view approach by partitioning the interaction space with two radii: a strong-connection radius $c_{1}$ and a weak-connection radius $c_{2}$. To explore the benefits of multi-view processing while maintaining a simple and interpretable relationship between the two scales, we enforce a fixed ratio by setting $c_{2}=2c_{1}$ in all our experiments. This allows the model to simultaneously capture immediate local neighborhoods ($\mathcal{G}_{1}$) and broader, second-order contextual regions ($\mathcal{G}_{2}$). The specific ranges for the systematic search over $c_{1}$ for each task, along with full implementation details, are provided in Appendix B.2.
@@ 7.2 Results and Analysis
A second key design parameter is the number of partitions in the graph-tuple. In this work we intentionally instantiate the graph-tuple with $k=2$ views corresponding to the strong and weak graphs. This two-scale design already captures the main separation between primary and contextual interactions in our dense graphs, while keeping the number of intra- and inter-scale message-passing paths manageable. Allowing more than two partitions would increase computational and tuning complexity by introducing additional operators and paths, but it could provide greater flexibility to capture diverse interactions. Beyond multi-scale analysis, our framework is also naturally suited to heterogeneous graphs, where each of the $k$ views may correspond to a distinct relation type. Systematically exploring these higher-order, multi-relational graph-tuples is another promising avenue for future research.
@@ A.2 Expressivity and Risk Dominance
These theoretical results directly apply to the linear backbone of the GNNs used in our experiments. Specifically, our analysis focuses on these linear operators and does not consider the nonlinear activation functions applied to their outputs. While a full characterization of the nonlinearities is more complex, our analysis provides a clean conceptual baseline: any nonlinear architecture built upon this backbone inherits the fundamental expressivity gap between the underlying operator classes. Extending such guarantees to fully nonlinear settings is an interesting but technically nontrivial direction that we leave for future work.
@@ B.1 Molecular Property Prediction
Since the QM7b dataset is feature-scarce, we first construct node and edge features from the molecule’s Coulomb matrix, $X$. Following prior work [blum2024learning], we derive initial features directly from the Coulomb matrix entries. Specifically, we apply a "binary expansion" technique to expand the scalar diagonal entries ($X_{ii}$) and off-diagonal entries ($X_{ij}$) into 100-dimensional vectors, which serve as the initial node and edge features, respectively. These raw scalar values are then projected into a 100-dimensional hidden space by learnable encoders.
@@
For training, all models use a batch size of 128. We train the models by minimizing the L1 Loss (Mean Absolute Error) using the Adam optimizer [kingma2014adam] with an initial learning rate of $5\times 10^{-3}$ and weight decay of $10^{-5}$. A cosine–plateau scheduler reduces the learning rate by a factor of 0.8 after five epochs without validation improvement (minimum $10^{-5}$). Early stopping is triggered after 20 idle epochs or when the run reaches a maximum of 1000 epochs.